### v1

# Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity

Abstract
We investigate the transition rates of a centripetally accelerated atom interacting with electromagnetic vacuum fluctuations inside a high-quality cavity. Our findings reveal that the emission and excitation rates can be extensively manipulated by adjusting the cavity’s normal mode frequency and the rotational angular velocity. Using experimentally feasible parameters, we demonstrate that, in one scenario, the excitation rate can reach magnitudes as high as $10^{7}~{}{\rm s}^{-1}$, while the emission rate remains negligible, trailing by 9 orders of magnitude. This suggests the potential for substantial population inversion in an ensemble of atoms. In another scenario, both the emission and excitation rates can simultaneously reach magnitudes as high as $10^{7}~{}{\rm s}^{-1}$, indicating that millions of transitions per second are expected even for a single atom. These remarkable results not only offer a new method for controlling the radiative properties of atoms but also open avenues for the experimental verification of the circular Unruh effect using cutting-edge quantum technologies.
Introduction.— The study of atomic transitions has long been central to our understanding of the interaction between matter and electromagnetic fields within the framework of quantum electrodynamics. In free space, an atom in its excited state can spontaneously emit a photon and transition to a lower energy state, a process known as spontaneous emission. Conversely, spontaneous excitation, where an atom in its ground state absorbs energy from the surrounding electromagnetic vacuum to transition to an excited state, is not observed for inertial atoms in free space. It is well understood that spontaneous emission in vacuum is governed by the interaction between the atom and the fluctuating electromagnetic field. Therefore, if vacuum fluctuations are somehow modified, changes in the atom’s radiative properties would naturally occur.
One particularly intriguing scenario arises when atoms are placed in a resonant cavity, as the presence of the cavity alters the density of states available for the fluctuating electromagnetic field. This modification can lead to significant changes in the transition rates of the atom, exemplified by the Purcell effect Purcell46 , first proposed in 1946, where the rate of spontaneous emission is enhanced due to the modified vacuum environment within the cavity. This enhancement was later observed in both the microwave Purcell2 and optical regimes Purcell3 . The Purcell effect has been instrumental in developing a deeper understanding of how confined fluctuating electromagnetic fields in vacuum can influence atomic behavior, opening up new avenues for manipulating and controlling quantum systems and light-matter interactions.
Vacuum fluctuations of the electromagnetic field can be profoundly altered not only by confinement to bounded spaces, such as cavities, but also by the non-inertial motion of atoms. Atoms in such motion interact with these modified vacuum fluctuations in unique ways, leading to phenomena that have no counterpart in inertial motion. For instance, an atom undergoing uniform linear acceleration perceives the vacuum, as seen by an inertial observer, as a thermal bath. This leads to the atom spontaneously transitioning to a higher energy level, a phenomenon known as the Unruh effect Davies75 ; Unruh76 ; Fulling73 . Additionally, the emission rate of an accelerated atom is also enhanced compared to that of an inertial atom, mimicking the behavior of an atom in a thermal environment. However, the excitation rate typically remains much lower than the emission rate, except under extremely high accelerations where it can approach the emission rate. Therefore, while the Unruh effect is theoretically compelling, it has remained elusive in direct experimental observation due to the extreme accelerations required.
At this point, it is important to note that the Unruh effect is significant not only for its intrinsic properties but also for its deep connection to Hawking radiation from black holes Hawking74 ; Hawking75 . Both phenomena are linked through the equivalence principle, which draws a connection between acceleration and gravity.
Similar to the case of linear acceleration, an observer undergoing uniform centripetal acceleration would also perceive the vacuum of an inertial atom as an environment with radiation, although the spectrum of this radiation is non-thermal Bell83 ; Hacyan1986 ; Bell87 ; Kim1987 ; Unruh98 ; Rosu2005 . Consequently, spontaneous excitation can occur for atoms under centripetal acceleration in a vacuum, a phenomenon that can be regarded as the circular Unruh effect. The emission rate in this scenario can also be enhanced by centripetal acceleration, although the behavior differs significantly from that under uniform linear acceleration Letaw1981 ; Takagi1984 ; Letaw1980 ; Rogers88 ; Davies1996 ; Lorenci2000 ; YaoJin2014-1 ; YaoJin2014-2 ; Biermann2020 ; Bunney23 ; Bunney23-2 ; yu24 .
Interestingly, recent studies have shown that when the rotational angular velocity exceeds the transition frequency of the atom, the excitation rate can reach the same order of magnitude as the emission rate, even when the orbital radius and thus the centripetal acceleration are extremely small yu24 . These findings not only challenge the conventional understanding that a significant Unruh effect requires a large acceleration and reveal that the radiative properties of a rotating atom can be profoundly different from those of inertial atoms, even at vanishingly small centripetal accelerations, but also naturally lead to further questions. Specifically, on one hand, how would the radiative properties of inertial atoms inside a cavity change if they are rotated? Can the maximum cavity-amplified emission rate for inertial atoms be further enhanced by centripetal acceleration, as it is in free space? On the other hand, how would the radiative properties of rotating atoms in free space change if they are placed inside a cavity? Can the excitation rate of non-inertial atoms be drastically amplified by cavities, similar to how the emission rate is for inertial atoms? If possible, can the excitation rate even exceed the emission rate, leading to population inversion? These questions are central to the present study, where we explore the interplay between the effects of atomic circular acceleration, a form of non-inertial motion, and the confinement of the electromagnetic field within the cavity, with an eye toward controlling atomic radiative properties and experimentally verifying the circular Unruh effect using cutting-edge quantum technologies.
Transition rates of centripetally accelerated atoms in a cavity.— We consider a two-level atom with a proper transition frequency $\omega_{0}$ coupled to fluctuating vacuum electromagnetic fields within a cavity. Assuming that the atom moves along a circular orbit, its trajectory in the laboratory frame is described by
$x(t)=R\cos(\Omega t),~{}y(t)=R\sin(\Omega t),~{}z(t)=0\;,$ (1)
where $R$ represents the radius of the orbit, $\Omega$ the rotational angular velocity, and $t$ the coordinate time. In the laboratory frame, the Hamiltonian characterizing the dipole interaction between the atom and the electromagnetic field can be expressed as
$H_{I}=-\sum_{m=\rho,\phi,z}D_{m}\mathcal{E}_{m}\;,$ (2)
where $D_{m}$ is the electric dipole moment operator of the atom, and the components of $\mathcal{E}_{m}$ are
$\mathcal{E_{\rho}}$ $=$ $\Omega RB_{z}+\cos(\Omega t)E_{x}+\sin(\Omega t)E_{y},$
$\mathcal{E_{\phi}}$ $=$ $[\cos(\Omega t)E_{y}-\sin(\Omega t)E_{x}]/\gamma,$
$\mathcal{E}_{z}$ $=$ $E_{z}-\Omega R\sin(\Omega t)B_{y}-\Omega R\cos(\Omega t)B_{x},$ (3)
with $\gamma=\left(1-{\Omega^{2}R^{2}}/{c^{2}}\right)^{-\frac{1}{2}}$ being the Lorentz factor, $c$ the speed of light, and $E_{i}$, $B_{i}$ $(i=x,y,z)$ the components of the electric and magnetic fields in the laboratory coordinate system, respectively.
Using perturbation theory, the emission rate $\Gamma_{\downarrow}$ and excitation rate $\Gamma_{\uparrow}$ of the atom in the laboratory frame can be derived as follows
[…]
where $t_{-}=t-t^{\prime}$, ${d}_{m}=\langle e|D_{m}|g\rangle$ is the dipole transition matrix element with $|g\rangle$ ($|e\rangle$) being the ground (excited) state of the atom, $G_{mn}(t_{-})=\langle 0|\mathcal{E}_{m}\left(t,\mathbf{x}\right)\mathcal{E}_{n}\left(t^{\prime},\mathbf{x}^{\prime}\right)|0\rangle$ are the two-point functions of the electromagnetic fields with $|0\rangle$ being the vacuum state of the electromagnetic fields, and $\omega=\omega_{0}(1-{\Omega^{2}R^{2}}/{c^{2}})^{\frac{1}{2}}$ is the transition frequency of the atom in the laboratory frame. The $+$ and -- signs in the exponential function correspond to the emission rate $\Gamma_{\downarrow}$ and excitation rate $\Gamma_{\uparrow}$, respectively. For simplicity, if we assume the transition matrix elements ${d}_{m}$ are real, then the cross terms in Eq. (4) vanish. For the explicit forms of the two-point functions $\langle 0|\mathcal{E}_{m}\left(t,\mathbf{x}\right)\mathcal{E}_{m}\left(t^{\prime},\mathbf{x}^{\prime}\right)|0\rangle$. These functions are dependent on the explicit form of the density of states within the cavity $\rho(\omega_{k})$, where $\omega_{k}$ denotes the frequency of the field mode. The density of states within a cavity is typically modeled as a Lorentzian distribution given by
$\rho(\omega_{k})=\frac{1}{\pi}\frac{\omega_{c}/Q}{(\omega_{c}/Q)^{2}+(\omega_{k}-\omega_{c})^{2}}\;,$ (5)
where $Q~{}(Q\gg 1)$ represents the quality factor, and $\omega_{c}$ denotes the normal mode frequency of the cavity. By substituting the explicit forms of the field correlation functions into Eq. (4), the leading terms of the emission rate $\Gamma_{\downarrow}$ in the non-relativistic limit, i.e., when the linear velocity $v=R\Omega$ is much smaller than the speed of light $c$, can be calculated as
$\Gamma_{\downarrow}\approx$ $\frac{\omega_{c}Q}{6\epsilon_{0}\hbar V}\bigg{\{}\frac{2\omega_{0}}{Q^{2}\left(\omega_{0}-\omega_{c}\right){}^{2}+\omega_{c}^{2}}d_{z}^{2}$
$+\left(d_{\rho}^{2}+d_{\phi}^{2}\right)\times\left[\frac{\omega_{0}+\Omega}{Q^{2}\left(\omega_{0}+\Omega-\omega_{c}\right){}^{2}+\omega_{c}^{2}}\right.$
$\left.+\frac{\left(\omega_{0}-\Omega\right)\Theta\left(\omega_{0}-\Omega\right)}{Q^{2}\left(\omega_{0}-\Omega-\omega_{c}\right){}^{2}+\omega_{c}^{2}}\right]\bigg{\}}.$ (6)
Meanwhile, the excitation rate $\Gamma_{\uparrow}$ is given by
$\Gamma_{\uparrow}\approx\frac{Q\omega_{c}\left(\Omega-\omega_{0}\right)\left(d_{\rho}^{2}+d_{\phi}^{2}\right)}{6\epsilon_{0}\hbar V\left[Q^{2}\left(\Omega-\omega_{0}-\omega_{c}\right){}^{2}+\omega_{c}^{2}\right]},$ (7)
for $\Omega>\omega_{0}$, and
$\Gamma_{\uparrow}\approx\frac{QR^{2}\omega_{c}\left(2\Omega-\omega_{0}\right){}^{3}\left(d_{\rho}^{2}+d_{\phi}^{2}\right)}{40c^{2}\epsilon_{0}\hbar V\left[Q^{2}\left(2\Omega-\omega_{0}-\omega_{c}\right){}^{2}+\omega_{c}^{2}\right]},$ (8)
for $\frac{\omega_{0}}{2}<\Omega\leq\omega_{0}$. Here $\epsilon_{0}$ is the vacuum permittivity, $V$ represents the volume of the cavity, $\Theta(x)$ is the Heaviside step function, which is equal to 00 when $x<0$ and equal to $1$ when $x>0$. Note that here we do not consider the case when $0<\Omega\leq\frac{\omega_{0}}{2}$, as the leading term of the excitation rate is of order $({R\Omega}/{c})^{4}$ or higher. In the following, we discuss explicitly the properties of the transition rates for centripetally accelerated atoms within a cavity, focusing on the interplay between the effects of centripetal acceleration and cavity. It should be noted that the analysis assumes $Q\gg 1$.
Rotational effect.— First, we examine the impact of rotation on the radiative properties of atoms by comparing the emission rate of centripetally accelerated atoms inside a cavity with that of inertial atoms in the same environment. In particular, we focus on whether the maximum cavity-amplified emission rate for inertial atoms can be further enhanced by centripetal acceleration.
When the atom is resonant with the cavity, i.e., when $\omega_{0}=\omega_{c}$, the emission rate for inertial atoms within the cavity reaches its maximum, given by the peak value:
$\Gamma_{\downarrow}^{\rm in}=\frac{Q}{3\epsilon_{0}\hbar V}\left(d_{z}^{2}+d_{\rho}^{2}+d_{\phi}^{2}\right).$ (9)
When the atoms are rotated, the emission rate becomes
$\Gamma_{\downarrow}\approx$ $\frac{Q}{3\epsilon_{0}\hbar V}d_{z}^{2}+\frac{Q\omega_{0}[\omega_{0}+\Omega+(\omega_{0}-\Omega)\,\Theta(\omega_{0}-\Omega)]}{6\epsilon_{0}\hbar V(Q^{2}\Omega^{2}+\omega_{0}^{2})}$
×$\times(d_{\rho}^{2}+d_{\phi}^{2}),$ (10)
where $Q\gg 1$. As the rotational angular velocity increases from zero, the contribution of the transverse polarization (atomic polarization perpendicular to the rotation axis) to the emission rate is strongly suppressed by rotation. Consequently, the primary contribution to the emission rate arises from the axial polarization (atomic polarization parallel to the rotation axis). According to Eq. (Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity), rotation always suppresses the emission rate, and the larger the rotational angular velocity, the smaller the emission rate becomes. This implies that, unlike in free space, the maximum cavity-amplified emission rate for inertial atoms cannot be further enhanced by centripetal acceleration when the atom is resonant with the cavity.
When the atom is not resonant with the cavity, the emission rate, as described by Eq. (Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity), initially increases with the rotational angular velocity. After reaching a peak, it decreases as the rotational speed continues to increase. The peak appears at $\Omega=\omega_{c}-\omega_{0}$ when $\omega_{0}<\omega_{c}$, and at $\Omega=\omega_{0}-\omega_{c}$ when $\omega_{0}>\omega_{c}$. In either case, the peak value of the emission rate is
$\Gamma_{\downarrow}\approx\frac{Q}{6\epsilon_{0}\hbar V}\left(d_{\rho}^{2}+d_{\phi}^{2}\right),$ (11)
where it is assumed that $|\omega_{0}-\omega_{c}|\gg\omega_{c}/Q$. In this scenario, the leading contribution to the emission rate comes from the transverse polarization. For inertial atoms in the same cavity, the emission rate is given by
$\Gamma_{\downarrow}^{\rm in}\approx\frac{1}{3\epsilon_{0}\hbar VQ}\frac{\omega_{0}\omega_{c}}{(\omega_{0}-\omega_{c})^{2}}\left(d_{z}^{2}+d_{\rho}^{2}+d_{\phi}^{2}\right).$ (12)
Given that $Q\gg 1$, the spontaneous emission is significantly enhanced by rotation if the atom’s transition frequency $\omega_{0}$ is not close to the cavity’s normal mode frequency $\omega_{c}$ (i.e., $|\omega_{0}-\omega_{c}|\gg\omega_{c}/Q$). This significant enhancement occurs despite the emission rate being dominated by the transverse polarization, which is markedly different from the behavior observed in inertial atoms. Both Eq. (Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity) and Eq. (11) are consistently smaller than Eq. (9), suggesting that the maximum cavity-amplified emission rate for inertial atoms cannot be further enhanced by rotation, regardless of whether the atom is resonant with the cavity.
Cavity effect.— Next, we investigate how the radiative properties of centripetally accelerated atoms can be significantly influenced by the presence of the cavity. To do this, we compare the peak values of the emission and excitation rates of these atoms when they are inside a cavity with those when they are in free space.
1. We start with the emission rate. As shown in Eq. (Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity), when the rotational angular velocity $\Omega$ exceeds the transition frequency of the atom $\omega_{0}$, i.e., when $\Omega>\omega_{0}$, the emission rate as a function of the cavity’s normal mode frequency exhibits two peaks, located at $\omega_{c}=\omega_{0}$ and $\omega_{c}=\omega_{0}+\Omega$, respectively. When $\Omega\leq\omega_{0}$, a third peak appears at $\omega_{c}=\omega_{0}-\Omega$. According to Eq. (Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity) and Eq. (11), for isotropically polarizable atoms with $d_{z}^{2}=d_{\rho}^{2}=d_{\phi}^{2}\equiv d^{2}$, the leading terms of the peak emission rates at these points take the same value, which is $\frac{Qd^{2}}{3\epsilon_{0}\hbar V}$. For comparison, the emission rate of a rotating atom in free space is yu24
$\Gamma_{\downarrow}^{\rm free}\approx$ $\frac{\omega_{0}^{3}}{3\pi\epsilon_{0}\hbar c^{3}}\left[\left(1+\frac{3\Omega^{2}}{\omega_{0}^{2}}\right)\left(d_{\rho}^{2}+d_{\phi}^{2}\right)+d_{z}^{2}\right],$ (13)
when $\Omega\leq\omega_{0}$, and
$\Gamma_{\downarrow}^{\rm free}\approx$ $\frac{\left(\Omega+\omega_{0}\right){}^{3}}{6\pi\epsilon_{0}\hbar c^{3}}\left[\left(d_{\rho}^{2}+d_{\phi}^{2}\right)+\frac{2\omega_{0}^{3}}{\left(\Omega+\omega_{0}\right){}^{3}}d_{z}^{2}\right],$ (14)
when $\Omega>\omega_{0}$. Comparing $\Gamma_{\downarrow}=\frac{Qd^{2}}{3\epsilon_{0}\hbar V}$ with Eq. (13) and Eq. (14) reveals that the emission rate of rotating atoms with isotropic polarizability can be enhanced by a factor of $\frac{Qc^{3}}{V\omega_{0}^{3}}$ when $\Omega\leq\omega_{0}$, and by a factor of $\frac{Qc^{3}}{V(\Omega+\omega_{0})^{3}}$ when $\Omega>\omega_{0}$, respectively. These enhancement factors can be extremely large, as $Q\gg 1$ and the cavity volume can be very small. As a result, the emission rate for rotating atoms inside the cavity is significantly amplified compared to that in free space yu24 .
2. Next, we examine whether the excitation rate of centripetally accelerated atoms can also be significantly amplified by the presence of the cavity. We begin with the scenario when $\Omega>\omega_{0}$. According to Eq. (7), the excitation rate peaks when the normal mode frequency of the cavity is tuned to $\omega_{c}=\Omega-\omega_{0}$, with the peak value given by
$\Gamma_{\uparrow}\approx\frac{Q}{6\epsilon_{0}\hbar V}\left(d_{\rho}^{2}+d_{\phi}^{2}\right),$ (15)
which coincides with the peak value of the emission rate Eq. (11) at $\omega_{c}=\omega_{0}+\Omega$, indicating that the excitation rate for rotating atoms inside the cavity can be drastically amplified, even reaching the maximum cavity-amplified emission rate for rotating atoms. For comparison, the excitation rate of a centripetally accelerated atom in free space when $\Omega>\omega_{0}$ is given by yu24 ,
$\Gamma_{\uparrow}^{\rm free}\approx\frac{(\Omega-\omega_{0})^{3}}{6\pi\epsilon_{0}\hbar c^{3}}\left(d_{\rho}^{2}+d_{\phi}^{2}\right),$ (16)
Therefore, the maximum excitation rate of a centripetally accelerated atom inside a cavity is approximately $\frac{Qc^{3}}{V(\Omega-\omega_{0})^{3}}$ times that in free space.
On the other hand, when $\frac{\omega_{0}}{2}<\Omega\leq\omega_{0}$, the excitation rate of rotating atoms within a cavity reaches its peak when the normal mode frequency of the cavity is tuned to $\omega_{c}={2\Omega-\omega}_{0}$ according to Eq. (8), with the peak value given by the following expression:
$\Gamma_{\uparrow}\approx\frac{QR^{2}\left(2\Omega-\omega_{0}\right){}^{2}}{40\epsilon_{0}\hbar Vc^{2}}\left(d_{\rho}^{2}+d_{\phi}^{2}\right).$ (17)
For comparison, the corresponding excitation rate for rotating atoms in free space is
$\Gamma_{\uparrow}^{\rm free}\approx\frac{R^{2}\left(2\Omega-\omega_{0}\right){}^{5}}{40\pi\epsilon_{0}\hbar c^{5}}\left(d_{\rho}^{2}+d_{\phi}^{2}\right).$ (18)
Comparing Eq. (17) with Eq. (18), it is evident that, since $Q\gg 1$, the presence of the cavity significantly amplifies the excitation rate of rotating atoms by a factor of $\frac{Qc^{3}}{V(2\Omega-\omega_{0})^{3}}$. This indicates that the excitation rate can be significantly amplified by a high-$Q$ cavity, regardless of whether the rotational angular velocity exceeds the atom’s transition frequency.
3. Generally, the peaks of the emission and excitation rates do not overlap. However, they can coincide when the rotational angular velocity takes specific values. For $\Omega>\omega_{0}$, the peak of the excitation rate at $\omega_{c}=\Omega-\omega_{0}$ coincides with the peak of the emission rate at $\omega_{c}=\omega_{0}$ when $\Omega=2\omega_{0}$. For isotropically polarizable atoms, the leading term of the emission rate Eq. (Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity) is the same as that of the excitation rate Eq. (15). This suggests that both the emission and excitation rates can be significantly enhanced simultaneously in a high-$Q$ cavity. For the case $\frac{\omega_{0}}{2}<\Omega\leq\omega_{0}$, when $\Omega=\omega_{0}$ and $\Omega=2\omega_{0}/3$, the peak of the excitation rate at $\omega_{c}=2\Omega-\omega_{0}$ coincides with the peak of the emission rate at $\omega_{c}=\omega_{0}$ and $\omega_{c}=\omega_{0}-\Omega$, respectively. In this scenario, the emission rate (given by Eq. (Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity) and Eq. (11)) is always larger than the excitation rate (given by Eq. (17)).
4. Now, we further investigate whether the cavity-enhanced excitation rate can surpass the emission rate. We begin with the case where $\Omega>\omega_{0}$. As discussed previously, this cannot occur when the peaks of the emission and excitation rates coincide. Thus, we focus on cases when the rotational angular velocity $\Omega$ is not in the vicinity of $2\omega_{0}$, thereby avoiding overlap between the peaks of the emission and excitation rates. According to Eq. (Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity), when $\Omega$ is not near $2\omega_{0}$, the emission rate for rotating atoms inside a cavity with the normal mode frequency tuned to $\omega_{c}=\Omega-\omega_{0}$ is given by
$\Gamma_{\downarrow}\approx$ $\frac{\Omega^{2}-\omega_{0}^{2}}{24\epsilon_{0}\hbar VQ\omega_{0}^{2}}\left[\left(d_{\rho}^{2}+d_{\phi}^{2}\right)+\frac{8\omega_{0}^{3}}{\left(\Omega-2\omega_{0}\right){}^{2}\left(\Omega+\omega_{0}\right)}d_{z}^{2}\right],$ (19)
where it is assumed that the quality factor is very large such that $Q\gg 1$ and $Q\gg\Omega/\omega_{0}$. Comparing Eq. (15) and Eq. (19), the excitation rate is approximately $\frac{Q^{2}}{(\Omega/\omega_{0})^{2}-1}$ times that of the emission rate when $\Omega\gtrsim 3\omega_{0}$. Remarkably, this suggests that the excitation rate can be significantly greater than the emission rate, given that $Q\gg 1$. Therefore, it is possible to achieve a significant population inversion for centripetally accelerated atoms within a high-$Q$ cavity, where the higher energy state becomes much more populated than the lower energy state.
The case when $\frac{\omega_{0}}{2}<\Omega\leq\omega_{0}$ can be analyzed similarly. When $\Omega$ is not close to $\Omega=\omega_{0}$ and $\Omega=2\omega_{0}/3$ such that the peaks of the emission and excitation rates overlap, the emission rate for rotating atoms inside a cavity with the normal mode frequency tuned to $\omega_{c}=2\Omega-\omega_{0}$ is given by
$\Gamma_{\downarrow}\approx$ $\frac{2\Omega-\omega_{0}}{6\epsilon_{0}\hbar VQ}\bigg{\{}\frac{2\left[4\Omega^{3}+\omega_{0}\left(\Omega^{2}-8\omega_{0}\Omega+4\omega_{0}^{2}\right)\right]}{(\Omega-2\omega_{0})^{2}(3\Omega-2\omega_{0})^{2}}$
×$\times\left(d_{\rho}^{2}+d_{\phi}^{2}\right)+\frac{\omega_{0}}{2\left(\Omega-\omega_{0}\right){}^{2}}d_{z}^{2}\bigg{\}}.$ (20)
According to Eq. (17) and Eq. (Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity), it is possible for the excitation rate to surpass the emission rate if the parameters are chosen appropriately.
Discussion.— We have the following comments based on the investigations presented above.
1. It is well known that in a vacuum, an inertial atom in its ground state cannot spontaneously transition to an excited state. However, this is possible for atoms that are uniformly accelerated. From an operational perspective, the spontaneous transition of an accelerated atom from the ground state to an excited state in a vacuum is a direct manifestation of the Unruh effect Davies75 ; Unruh76 ; Fulling73 . The Unruh effect is fundamentally significant but has proven elusive in experimental detection, primarily because the spontaneous excitation is too faint at the accelerations achievable in current experiments. Although it has been shown in Ref. yu24 that the excitation rate of a rotating atom in a vacuum can reach the same order of magnitude as the emission rate when the rotational angular velocity exceeds the atomic transition frequency, resulting in a much higher excitation rate than that of an atom under uniform linear acceleration, the actual excitation rate remains extremely weak, making it difficult to observe in a laboratory setting. To illustrate this explicitly, we take $d=10^{-29}\,\text{Cm}$, $V=10^{-14}\,\text{m}^{3}$, $Q=10^{7}$, $R=50\,\text{nm}$, $\Omega=5\,\text{GHz}$, and $\omega_{0}=10\,\text{MHz}$, satisfying $R\Omega/c\ll 1$. Note that these parameters are experimentally feasible, and consistent with Ref. Lochan20 . With these values, the excitation rate in free space is found to be only $10^{-11}\,{\text{s}}^{-1}$, which is far below the threshold for experimental detection. In stark contrast, if the rotating atom is placed inside a cavity with the normal mode frequency tuned to $\omega_{c}=\Omega-\omega_{0}$, the excitation rate can be amplified to $10^{7}~{}{\text{s}}^{-1}$ according to Eq. (15), which exceeds that in free space by 18 orders of magnitude. Remarkably, this peak value is independent of the atom’s transition frequency and rotational speed, making the choice of parameters less restrictive. The only requirement is the condition $\omega_{c}=\Omega-\omega_{0}$. Meanwhile, the emission rate in this case is negligible, trailing the excitation rate by 9 orders of magnitude. This result highlights the potential for substantial population inversion. Such a dramatic enhancement of the excitation rate could facilitate the experimental detection of the circular Unruh effect and make it a highly feasible goal with current technology.
2. In addition to measuring the cavity-enhanced spontaneous excitation as previously discussed, the circular Unruh effect can also be verified by detecting cavity-enhanced spontaneous emission. By tuning the normal mode frequency of the cavity to $\omega_{c}=\Omega+\omega_{0}$ and using experimentally feasible parameters as before, it is possible to observe emission specifically due to centripetal acceleration, which can reach as high as $10^{7}\,{\text{s}}^{-1}$. More interestingly, by setting the rotational angular velocity and the cavity’s normal mode frequency to $\Omega=2\omega_{c}=2\omega_{0}=5\,\text{GHz}$ while keeping the other parameters unchanged, both the emission and excitation rates can reach as high as $\sim 10^{7}\,\text{s}^{-1}$ simultaneously. This configuration can result in millions of transition events per second for a single centripetally accelerated atom, offering a robust method for experimentally verifying the circular Unruh effect. Such a phenomenon is impossible for an inertial atom, even if its emission rate is higher when $\omega_{c}=\omega_{0}$, because once it emits a photon and transitions to the ground state, it cannot be excited again. In contrast, the continuous transitions between the ground and excited states of a single centripetally accelerated atom serve as compelling evidence for the circular Unruh effect.
3. Based on the discussions above, we argue that manipulating the transition rates and achieving significant population inversion are indeed feasible with current state-of-the-art technologies. Recently, hyperfast rotation of an optically levitated nanoparticle has been achieved by transferring the spin angular momentum of light to the particle’s mechanical angular momentum Reimann2018 ; Li2018 ; Zhang2021 . This progress opens up the promising possibility of attaching an atom to such a rapidly rotating, optically levitated nanoparticle. Notably, the physical parameters used in our numerical investigations ($R=50\,\text{nm}$ and $\Omega=5\,\text{GHz}$) align well with recent experimental developments Zhang2021 , suggesting that the proposed experiments could be practically realizable.
4. Finally, we note that there are recent studies focusing on the transition rates of accelerated atoms within a cavity Scully03 ; Lochan22 ; Lochan20 ; guo24 . Notably, in a recent work Lochan20 where only the emission of excited atoms is investigated, it was demonstrated that the emission rate of a rotating atom due to centripetal acceleration within a cavity can be significantly amplified by the cavity. However, the emission rate reported in Ref. Lochan20 is substantially lower–by 14 orders of magnitude–than the rate found in the current study when using the same parameters. This significant discrepancy arises because, in Ref. Lochan20 , the atoms are assumed to be polarizable only along the direction orthogonal to the plane of rotation. In contrast, our study assumes isotropic polarizability, which is more common in nature. In the non-relativistic limit ($R\Omega/c\ll 1$), the leading term for the emission rate when $\omega_{c}=\Omega+\omega_{0}$ is given by Eq. (11), where the dominant contribution comes from transverse polarization. This rate is approximately $\sim c^{2}/(R\Omega)^{2}~{}(\gg 1)$ times that from axial polarization, making the axial contribution calculated in Ref. Lochan20 negligible.
Summary.— We have investigated the transition rates of a centripetally accelerated atom coupled with fluctuating electromagnetic fields in a vacuum inside a resonant cavity. Our study reveals that the interplay between atomic centripetal acceleration, a form of non-inertial motion, and the confinement of the electromagnetic fields within a high-$Q$ cavity can profoundly alter the radiative properties of atoms. While the maximum cavity-amplified emission rate for inertial atoms cannot be further enhanced by rotation, the spontaneous emission can still be significantly amplified if the atom is far from resonance with the cavity. Notably, the transition rates change dramatically when rotating atoms are placed inside the cavity, enabling significant enhancement of either the emission or excitation rates, individually or even simultaneously.
Remarkably, with realistic physical parameters, we show that, in one scenario, the excitation rate can reach values as high as $10^{7}~{}{\rm s}^{-1}$, surpassing those in free space by 18 orders of magnitude. Meanwhile, the emission rate is negligible compared with the excitation rate, trailing by 9 orders of magnitude, which suggests the potential for significant population inversion in an ensemble of atoms. In another scenario, the emission and excitation rates can simultaneously reach magnitudes as high as $10^{7}~{}{\rm s}^{-1}$, indicating that millions of transitions per second are possible even for a single atom.
These findings not only offer a novel approach to controlling the radiative properties of atoms but also pave the way for the experimental verification of the circular Unruh effect using cutting-edge quantum technologies.
Acknowledgements.
This work was supported in part by the NSFC under Grant No. 12075084, No. 12375047, the innovative research group of Hunan Province under Grant No. 2024JJ1006, the Hunan Provincial Natural Science Foundation of China under Grant No. 2023JJ40515, and the Scientific Research Program of Education Department of Hunan Province of China under Grant No. 22B0762.